David's novel has 12 chapters. Some of the chapters have 25 pages and some of the chapters have 37 pages. How many of each size chapter does the novel have if it is 360 pages long?
Answer by richwmiller(17219)

a+b=12,
25a+37b=360
or
The chapters with 37 pages must come in groups of 5 to end in 5
37*5=185
360-185=175
175/25=7
5 chapters of 37 pages and 7 chapters of 25 pages